1. Using the FINRA Bond Center search engine, search for SEC 144A private placement corporate bonds. Using the advanced search engine, set sub-product type to “Corporate Bond”, check the box to “include” 144A Bonds, set the Moody’s rating from C to Caa, the Standard and Poor’s rating from D to BBB-, and the Trace Grade to “High Yield”. Click on SHOW RESULTS and sort from highest to lowest maturity. Display the details for the bond issued by “Chesapeake Energy Corp”. Make a screen capture of at least the first 10 lines of the display, and include it below.

2. A 30-year 6% semi-annual coupon bond has a tenor of 14 years and a yield to maturity of 7.000%. If the PAR value of the bond is $1000, what is the price of the bond today?
Answer #1

finra-markets.morningstar.com/BondCenter/Results.jsp Bonds EDIT SEARCH Bond Results Coupon Maturity 8.000 05/01/2022 7.000 11finra-markets.morningstar.com/BondCenter/Results.jsp CUIT SCARCA Bond Results Coupon Maturity 6.875 08/01/2025 Ratings Moody

finra-markets.morningstar.com/BondCenter/Results.jsp Coupon Maturity 10.750 12/15/2023 6.125 08/15/2025 Ratings Moodys S&P C

finra-markets.morningstar.com/BondCenter/Results.jsp Coupon Maturity 9.250 10/15/2024 Ratings Moodys® S&P Caa2 B- Last Sale

finra-markets.morningstar.com/BondCenter/Results.jsp Coupon Maturity 9.375 05/01/2024 Ratings Moodys® S&P Ca C Last Sale Pri

A finra-markets.morningstar.com/BondCenter/Results.jsp Coupon 7.500 Ratings Maturity Moodys S&P 08/01/2021 Caa3 CCC Last Sal

• finra-markets.morningstar.com/BondCenter/Results.jsp Coupon Maturity 7.375 10/15/2024 7.250 10/15/2025 Ratings Moodys® S&P

finra-markets.morningstar.com/BondCenter/Results.jsp Coupon 9.125 Maturity 06/30/2025 Ratings Moodys® S&P Caa1 BB- Last Sale

5 + 1 tinra-markets.morningstar.com/BondCenter/BondDetail.jsp?ticker=FCHR4499459&symbol=CHR4499459 Market Data Investors > Ma

finra-markets.morningstar.com/BondCenter/BondDetail.jsp?ticker=FCHK4499459&symbol=CHK4499459 Price/Yield Chart Classification

2.

Coupon Rate = 6% for annual = 3% for semi-annual bond

Coupon value = 3% * 1000 = $ 30

Tenor = Maturity = 14 years = 28 years for semi-annual

YTM = 7% = 3.5% for semi-annual

Price of the bond = 30/ ((1+0.035) ^ 1) + 30/ ((1+0.035) ^ 2) + 30/ ((1+0.035) ^ 3) + 30/ ((1+0.035) ^ 4) + ---------------------------------------- Likewise + 30/ ((1+0.035) ^ 28) + 100/ ((1+0.035) ^ 28) ( Coupon value are discounted at YTM + Par value received at maturity discounted at YTM)

Right hand side formed a geometric progression with first term = 30/1.035 = 28.99, ratio,r = 1/1.035 = 0.97, n=28

Price of the bond = 28.99 * ( 1 – ((0.97) ^ 28) )/ ( 1 – 0.97) + 1000 / ( 1.035 ^ 28)

Price of the bond = $ 911.67